Water vapor initially at 10 bara and 400°C is contained within a piston-cylinder assembly. The water is cooled at constant volume until its temperature is 150°C. The water is then condensed isothermally to saturated liquid. Sketch the process in T-v coordinates. Clearly label the three states. For the water as the system, (a) write the symbolic equation for the specific work (and simplify), w; and (b) evaluate the specific work, in kJ/kg.
Answer #1

solution P= lotar T = 400 a . . water state L it is given that P = 10 bar = 1 MPa T= 400c from steam table [ CATT-] U; 2..306for this we have 7 = 150°C = ∞ = 0.00109 m2/ky & P = b = 0.4749 MPa 00103 -3064 vu (a) work for process 1-2 Wire - sear since(b) specific work hazz POV = P(ų) E 0.4759x10^x [00109 - 0.30663 = -0.145392 X106 Illy - - 145.392X16 = - 145.392 KJlk * Tot